Circular No. 6917

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998cg IN ANONYMOUS GALAXY A. Gal-Yam and D. Maoz, Tel Aviv University, report their discovery

of an apparent supernova (R ∼ 18.5) on unfiltered CCD images taken with the Wise Observatory 1-m telescope on May 1. The supernova is located at α = 12^{h}18^{m}18^{s}.13, δ = +20^{o}44′26′′.05 (equinox 2000.0), which is 0′′.4 east and 4′′ south of the nucleus of its host galaxy in the foreground of the galaxy cluster Abell 1514. The star was not visible in images obtained on Apr. 6 (limiting mag R ∼ 21), but was confirmed on CCD images taken by J. Dan on May 13 (R ∼ 19.5) with the same telescope. Further details and finding charts can be obtained via ftp://wise3.tau.ac.il/ftp/

pub/dani/SN1.ps or by e-mail (avishay@wise.tau.ac.il).

SUPERNOVA 1998ch IN ANONYMOUS GALAXY L. Germany reports the discovery of an apparent supernova (V = 20.3)

by the Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC 6639, 6893) on V and R CCD images taken on May 18 (limiting mag V = 20.5) by S. Chan at the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1998ch is located near Abell 3559 at α = 13^{h}29^{m}13^{s}.31, δ = −29^{o}17′26′′.8 (equinox 2000.0), and is offset 1′′.3 west and 1′′.8 south from the center of its host galaxy. A nearby star has position end figures 14^{s}.53, 16′52′′.6. SN 1998ch was not visible on similar images taken on Apr. 30 (limiting mag V = 21.5) but was confirmed by S. Sabine with the same telescope on May 26 (limiting mag V = 21.5).

SUPERNOVA 1998ce IN MCG –4-24-19 J. Maza, Department of Astronomy, University of Chile, reports the

independent discovery of this supernova (B ∼ 16.5) by Marina Wischnjewsky, on a T-Max 400 film taken by L. E. González with the Maksutov telescope at Cerro El Roble on May 19.99 UT. The object was not present on a film obtained on Apr. 29 (B > 19).

SUPERNOVA 1997eg IN NGC 5012 J. Mueller reports that a IV-N plate taken on May 16 with the 1.2-m

Oschin Telescope in the course of the second Palomar Sky Survey shows SN 1997eg at mag 16.

1998 May 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green